4 d2y/dx2+25y=0


1
Expert's answer
2021-09-30T09:28:45-0400

Characteristic equation: "4r^2+25=0"

"r=\frac{5}{2} i, i=-\frac{5}{2}i".

General solution of DE: "y=C_1sin(\frac{5}{2}x)+C_2cos(\frac{5}{2}x)".